# Quantum Pufferfish Privacy: A Flexible Privacy Framework for Quantum Systems

### v1

@@ VI-B Analysis of Depolarization Mechanism
In this section, we instantiate $\mathcal{A}$ to be the depolarizing channel with parameter $p$, denoted as $\mathcal{A}^{p}_{\mathrm{Dep}}$ and defined previously in (127).
@@
$\mathsf{U}(\mathcal{A}_{\mathrm{Dep}}^{p})=\frac{p(d^{2}-1)}{d^{2}}\leq\gamma.$ (168)
@@
Our proof of Proposition 7 relies on identifying that the utility metric is minimized by setting $\mathcal{B}=\mathcal{I}$, and then evaluating $\left\|\mathcal{I}-\mathcal{A}_{\mathrm{Dep}}^{p}\right\|_{\diamond}$ using the Choi states of the channels $\mathcal{A}_{\mathrm{Dep}}^{p}$ and $\mathcal{I}$, due to the joint covariance of the two channels under unitaries (i.e., $\mathcal{A}_{\mathrm{Dep}}^{p}\circ\mathcal{U}=\mathcal{U}\circ\mathcal{A}_{\mathrm{Dep}}^{p}$ and $\mathcal{I}\circ\mathcal{U}=\mathcal{U}\circ\mathcal{I}$ for every unitary channel $\mathcal{U}$).
@@
where: (a) follows from the unitary invariance of the diamond norm with $\mathcal{U}$ representing a unitary channel; (b) from the commutative property of $\mathcal{A}_{\mathrm{Dep}}^{p}$ with every unitary channel; (c) from the left-hand side being independent of $\mathcal{U}$; and (d) from the convexity of the diamond norm.
Next, observe that $\mathcal{B}^{\star}\coloneqq\int\,\mathsf{d}\mathcal{U}\ \mathcal{U}\circ\mathcal{B}\circ\mathcal{U}^{\dagger}$ is a quantum channel, and this is equal to a depolarization channel [50]. Then, $\mathcal{B}^{\star}=\mathcal{A}_{\mathrm{Dep}}^{q}$ for some $q\in[0,1]$.
@@
With the property of joint covariance of $\mathcal{I}$ and $\mathcal{A}_{\mathrm{Dep}}^{p}$ under unitaries [27, Proposition 7.82], we simplify this to
@@
Remark 10 (Employing other metrics of utility: fidelity and trace distance).
Employing trace distance as the metric, we define $\beta$-utility as
$\sup_{\rho\in\mathcal{D}(\mathcal{H})}\frac{1}{2}\left\|\rho-\mathcal{A}(\rho)\right\|_{1}\leq\beta,$ (181)
where $\mathcal{D}(\mathcal{H})$ is the set of all density matrices. Then, for the depolarization mechanism we have
$\sup_{\rho\in\mathcal{D}(\mathcal{H})}\frac{1}{2}\left\|\rho-\mathcal{A}_{\mathrm{Dep}}^{p}(\rho)\right\|_{1}=\frac{p(d-1)}{d}\leq\beta,$ (182)
which follows by observing that the supremum is achieved on pure states (due to convexity of the trace norm) and by following arguments similar to those in the proof of Proposition 7. Observe that, similar to Proposition 7, the utility degrades linearly with $p$ (recall that utility degrades with increasing $\beta,\gamma$).
Alternatively, we can consider the following utility metric based on fidelity:
$\inf_{\rho\in\mathcal{D}(\mathcal{H})}\mathsf{F}\!\left(\rho,\mathcal{A}(\rho)\right)\leq\zeta.$ (183)
Compared to the other metrics, by fidelity being a similarity metric rather than a distinguishability metric, higher $\zeta$ implies higher utility gain. For the depolarization mechanism, we have
$\inf_{\rho\in\mathcal{D}(\mathcal{H})}\mathsf{F}\!\left(\rho,\mathcal{A}^{p}_{\mathrm{Dep}}(\rho)\right)=1-p+\frac{p}{d}\leq\zeta.$ (184)
This follows due to the joint concavity of the root fidelity, so that the infimum is achieved by pure states. Lastly, consider that
$\mathsf{F}\left(\mathcal{A}_{\mathrm{Dep}^{p}}(\psi),\psi\right)$ $=\langle\psi|\mathcal{A}^{p}_{\mathrm{Dep}}(\psi)|\psi\rangle$ (185)
$=\langle\psi|\left[(1-p)|\psi\rangle\!\langle\psi|+\frac{\mathrm{I}_{d}}{d}\right]|\psi\rangle$ (186)
$=(1-p)|\langle\psi|\psi\rangle|^{2}+\frac{p}{d}\langle\psi|\psi\rangle$ (187)
$=(1-p)+\frac{p}{d},$ (188)
to arrive at the desired expression.
Next, we focus on understanding the privacy-utility tradeoff in terms of the parameter $p$. This can also be understood as figuring out whether a given noisy quantum channel, represented as a depolarization mechanism, is suitable for a given application when privacy and utility gains are predetermined.
Figure 6: For fixed $d=2$, the figure depicts the optimum utility changing with parameter $p$ and the achievability of $\varepsilon$-QPP changing with parameter $p$ (any $\varepsilon,\gamma$ values above the drawn curves are achievable for a fixed $p$ value). The value of $K$ encodes the domain knowledge available, where $K=1$ corresponds to no such additional information being available.
From Proposition 7, to achieve $\gamma$-utility, we require that
$p\leq\frac{\gamma d^{2}}{(d^{2}-1)}.$ (189)
In terms of privacy, to achieve $\varepsilon$-QPP in the chosen privacy framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$, from Theorem 3, we require that
@@
These two inequalities showcase the privacy-utility tradeoff of the depolarization mechanism. Thus, it is important to tune $p$ based on the required utility characterized by $\gamma$, as well as the privacy parameter $\varepsilon$.
Furthermore, from Theorem 3, we can also deduce that for a fixed $p$ value, the achievable privacy regime includes
$\varepsilon\geq\ln\!\left(1+\frac{(1-p)dK}{p}\right).$ (191)
Fig. 6 depicts the privacy-utility tradeoff in terms of the parameter $p$. Lower $p$ values achieve lower privacy (higher $\varepsilon$) while the utility gain is higher (lower $\gamma$). Privacy is relatively more sensitive to $p$ compared to utility, which degrades linearly with $p$. We also observe that with the reduction of the value $K$, the achievable privacy regime grows. The parameter $K$ from Theorem 3 encodes the domain knowledge available, which is fed into the privacy model from the $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ QPP framework. This also highlights the importance of encoding such information when available to achieve higher utility gains while providing the required privacy guarantees.
Next, we emphasize the impact of the dimension on the achievability of privacy and utility. In Fig. 7, we observe that the privacy level is more sensitive to the dimension than the utility. Considering the utility of the depolarization mechanism (given by $\frac{p(d^{2}-1)}{d^{2}}$), we can always have the following upper bound on it for every $d$:
$\frac{p(d^{2}-1)}{d^{2}}\leq p,$ (192)
where the upper bound is achieved when $d\to\infty$. However, the achievable privacy level in (191) degrades no larger than order $\ln(d)$. Thus, it is important to identify the right/optimum privacy parameters, especially in high-dimensional setups.
Figure 7: For fixed $K=1$, the figure depicts the optimum utility changing with parameter $p$ and the achievability of $\varepsilon$-QPP changing with parameter $p$ (any $\varepsilon,\gamma$ values above the drawn curves are achievable for a fixed $p$ value) for $d\in\{2,4,8\}$.
In the above analysis for the depolarization mechanism in Fig. 4, we chose $\mathcal{E}=\mathcal{I}$. It is an interesting future work to analyze utility for user-specific $\mathcal{E}$ channels: i.e., choosing $\inf_{\mathcal{B}\in\mathrm{CPTP}}\|\mathcal{E}-\mathcal{B}\circ\mathcal{A}^{p}_{\mathrm{Dep}}\circ\mathcal{E}\|_{\diamond}$ as the utility metric. If $\mathcal{E}$ possesses certain symmetries, then one can make use of arguments similar to those presented previously.
VI-C Characterizing Optimal Privacy-Utility Tradeoff when $\mathcal{Q}$ is a Small Set
In this subsection, we focus on identifying the optimal utility that can be obtained by applying an $(\varepsilon,\delta)$-QPP mechanism. Here, we first focus on the setting in which $\mathcal{Q}=\{(\mathcal{R}_{1},\mathcal{R}_{2}),(\mathcal{R}_{2},\mathcal{R}_{1})\}$, ¯=$\bar{\mathcal{M}}=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}$, and $\Theta=\{P_{X}\}$, but the following ideas can be extended to the case when $\mathcal{Q}$ is an arbitrary finite set and $\Theta$ includes finite number of probability distributions. However, the computational complexity involved in identifying the optimal utility increases with the cardinality of the set $\mathcal{Q}$ and $\Theta$, due to the addition of more constraints to the optimization problem.
To incorporate privacy requirements, we use the equivalent formulation of QPP via the DL divergence presented in Proposition 1. To this end, first, we employ the SDP formulated in Lemma 1 to compute the relevant DL divergence and then use that in the optimization of utility. We showcase the use of this SDP in characterizing optimal utility next. It is also important to note that this SDP formulation may also be used as a way to quantify privacy in a given $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$ framework, as presented in Proposition 1.
Proposition 8 (Optimal utility for fixed privacy constraints).
$\mathsf{U}\!\left(\varepsilon,\delta,\mathcal{R}_{1},\mathcal{R}_{2}\right)\coloneqq\\ \inf_{\begin{subarray}{c}\mu\geq 0\\ Z_{AD}\geq 0\\ \Gamma_{CD}^{\mathcal{B}}\geq 0\\ \Gamma_{AC}^{\mathcal{A}}\geq 0\\ \lambda_{1}\geq 0,Y_{1}\geq 0\\ \lambda_{2}\geq 0,Y_{2}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\mu:\\ Z_{AD}\geq\Gamma_{AD}-\mathrm{Tr}_{C}\!\left[\Gamma_{CD}^{\mathcal{B}}\mathrm{T}_{C}(\Gamma_{AC}^{\mathcal{A}})\right],\\ \mu I_{A}\geq\mathrm{Tr}_{D}\!\left[Z_{AD}\right],\\ \mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{C},\\ \mathrm{Tr}_{D}\!\left[\Gamma_{AC}^{\mathcal{A}}\right]=I_{A},\\ \ln(\lambda_{1})\leq\varepsilon,\\ \mathrm{Tr}\!\left[Y_{1}\right]\leq\delta,\\ Y_{1}\geq\mathrm{Tr}_{A}\!\left[\big{(}\mathrm{T}(\rho^{\mathcal{R}_{1}})\otimes I_{C}\big{)}\Gamma_{AC}^{\mathcal{A}}\right]\\ \quad-\lambda_{1}\mathrm{Tr}_{A}\!\left[(\mathrm{T}(\rho^{\mathcal{R}_{2}})\otimes I_{C})\Gamma_{AC}^{\mathcal{A}}\right],\\ \ln(\lambda_{2})\leq\varepsilon,\\ \mathrm{Tr}\!\left[Y_{2}\right]\leq\delta,\\ Y_{2}\geq\mathrm{Tr}_{A}\!\left[(\mathrm{T}\big{(}\rho^{\mathcal{R}_{2}}\big{)}\otimes I_{C})\Gamma_{AC}^{\mathcal{A}}\right]\\ \quad-\lambda_{2}\mathrm{Tr}_{A}\!\left[\big{(}\mathrm{T}(\rho^{\mathcal{R}_{1}})\otimes I_{C}\big{)}\Gamma_{AC}^{\mathcal{A}}\right]\end{array}\right\}.$ (193)
Proof:
The proof follows from the SDP formulation of the $\gamma$-utility given in Proposition 6, and the privacy constraints (i.e., $\max\{\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}_{1}})\|\mathcal{A}(\rho^{\mathcal{R}_{2}})\right),\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho^{\mathcal{R}_{2}})\|\mathcal{A}(\rho^{\mathcal{R}_{1}})\right)\}\leq\varepsilon$) imposed through the SDP formulation of DL divergence presented in Lemma 1. We also used the fact that for a superoperator $\mathcal{A}$ from system $A$ to $C$, the following equality holds [27, Eq. (3.2.14)]
This concludes the proof. ∎
can also be incorporated to compute the optimum utility.
Remark 12 (Optimal privacy parameters for fixed utility).
$\lambda^{\star}_{1}(\mathcal{A},\gamma,\delta)\coloneqq\\ \inf_{\begin{subarray}{c}\lambda\geq 0\\ Z_{AD}\geq 0\\ \Gamma_{CD}^{\mathcal{B}}\geq 0\\ Y_{1}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\lambda:\\ Z_{AD}\geq\Gamma_{AD}-\mathrm{Tr}_{C}\!\left[\Gamma_{CD}^{\mathcal{B}}\mathrm{T}_{C}(\Gamma_{AC}^{\mathcal{A}})\right],\\ \gamma I_{A}\geq\mathrm{Tr}_{D}\!\left[Z_{AD}\right],\\ \mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{C},\\ \mathrm{Tr}\!\left[Y_{1}\right]\leq\delta,\\ Y_{1}\geq\mathrm{Tr}_{A}\!\left[\big{(}\mathrm{T}(\rho^{\mathcal{R}_{1}})\otimes I_{C}\big{)}\Gamma_{AC}^{\mathcal{A}}\right]\\ \quad-\lambda\mathrm{Tr}_{A}\!\left[(\mathrm{T}(\rho^{\mathcal{R}_{2}})\otimes I_{C})\Gamma_{AC}^{\mathcal{A}}\right]\end{array}\right\}.$ (196)
$\varepsilon^{\star}(\mathcal{A},\gamma,\delta)\coloneqq\ln\!\left(\max\{\lambda^{\star}_{1}(\mathcal{A},\gamma,\delta),\lambda^{\star}_{2}(\mathcal{A},\gamma,\delta)\right)\}.$ (197)
@@ VII Auditing Privacy Frameworks
Auditing for privacy aims to detect violations in privacy guarantees and reject incorrect algorithms (see [51, 52, 53, 54] for classical approaches). In this section, we focus on how we can employ the tools of quantum information theory, as well as quantum algorithms, to audit the privacy of quantum systems. Here we focus on auditing algorithms for QDP guarantees, and note that the following ideas can be extended to auditing algorithms for privacy guarantees demanded by QPP.
The main idea behind auditing classical privacy frameworks is to translate the privacy requirement to a weaker privacy notion that is efficiently computable. Then, for algorithms that do not satisfy the privacy demanded by the relaxed privacy notion, we can conclude that the original privacy requirement is also violated. The pitfall of this approach is the impossibility of quantifying the gap between the original and relaxed privacy notions. For instance, knowing that the relaxed privacy notion is satisfied does not determine whether the original privacy requirement is also satisfied. However, here we target auditing QDP without translating it to a relaxed privacy notion. Note that the following ideas can be extended to auditing for a relaxed privacy notion as well (for the variants of QPP in Section IX).
@@ VII-A Techniques for Auditing QDP
Using the equivalent formulation in Proposition 1 and combining with the specific choices of $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ as given in Section III-C1, for $(\varepsilon,\delta)$-QDP, we have that
@@
Then, we can compute the left-hand side above by using the SDP formulation of $\overline{\mathsf{D}}^{\delta}\!\left(\cdot\|\cdot\right)$ presented in Lemma 1. This approach is helpful for low-dimensional setups because the time complexity of an SDP computation is polynomial in the dimension of the quantum states. However, the complexity of this approach grows exponentially with the number of qubits. Developing techniques to evaluate the DL divergence using a quantum computer would lead to methods for auditing procedures as well, which is beyond the scope of this work and thus left as a future work.
Next, using the equivalent formulation in Remark 5, consider that $(\varepsilon,\delta)$-QDP is equivalent to
@@
with $\gamma\geq 1$ for quantum states $\rho$ and $\sigma$ [10, Eq. (II.2)]. As shown in (43) and (45), the quantity on the right-hand side of (200) can be evaluated by means of an SDP, which implies the same for the hockey-stick divergence $\mathsf{E}_{\gamma}(\rho\|\sigma)$. Then, auditing QDP reduces to computing $\mathsf{E}_{\gamma}(\rho\|\sigma)$ for $\rho\sim\sigma$. However, as before, the time complexity of this SDP grows exponentially in the number of qubits. Based on the results of [55, 56] and the connection of $\mathsf{E}_{\gamma}(\rho\|\sigma)$ to the trace distance given in (201), we expect that evaluating $\mathsf{E}_{\gamma}(\rho\|\sigma)$ should be computationally difficult even for quantum computers [55, 56]. Nevertheless, there are proposals for evaluating the trace distance using variational quantum algorithms [57, 58] (which however do not give particular runtimes), and for cases in which the quantum states have low rank [59]. In what follows, we consider one approach from [57, 58] (via variational algorithms with parameterized quantum circuits) to estimate the quantity $\left\|\mathcal{A}(\rho)-e^{\varepsilon}\mathcal{A}(\sigma)\right\|_{1}$. Using such an estimate, for a fixed value of the privacy parameter $\varepsilon$, we can validate on which values of $\delta$ the needed guarantees are satisfied.
VII-A1 Using Quantum Circuits to Audit QDP
@@
Firstly, let us focus on how to estimate $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)$ for a fixed $\rho\sim\sigma$ and $\varepsilon$. With the ideas developed in [57, Algorithm 14], we discuss how a so-called quantum interactive proof can be used for estimating the privacy level. For that, refer to the quantum circuit in Fig. 8: the unitaries $U^{\rho}$ and $U^{\sigma}$ are used to prepare the states $\rho$ and $\sigma$, by tracing out systems $R_{1}$ and $R_{2}$, respectively. Then the algorithm $\mathcal{A}$ is applied on the systems $S_{i}$ for $i\in\{1,2\}$. The unitary $Q$ takes inputs $F_{i},B_{i}$ and outputs $F_{i}^{\prime},T_{i}$, where $F_{i}$ and $T_{i}$ are qubit systems. Finally, both of the $T_{i}$ systems are measured in the standard basis $\{|0\rangle,|1\rangle\}$, and the (classical) output random variable is denoted as $X_{\rho}$ for $i=1$ and $X_{\sigma}$ for $i=2$. Here $X_{\rho}$ and $X_{\sigma}$ take values in $\{0,1\}$. This procedure is repeated a sufficient number of times, and we use the results to estimate $\mathbb{P}(X_{\rho}=0)$ and $\mathbb{P}(X_{\sigma}=1)$.
@@
In quantum complexity terminology, this action could be conducted by a quantum prover who has unbounded computational power (we discuss how to relax this assumption in Remark 13). From [57, Eq. (128)] and the discussion therein, we conclude that
@@
Finding the $Q$ which achieves the optimum utility in (202) is practically infeasible. To relax this assumption, we replace the action of the prover (who finds this $Q$) with a parameterized circuit $Q_{\theta}$. Then we can use (202) as a utility function for training a variational quantum algorithm [60, 61] to estimate a lower bound for (203). With that we obtain a lower bound on $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)$, which we denote as $\mathsf{E}^{\mathrm{LB}}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)$. A lower bound gives a sufficient condition for ruling out algorithms that do not satisfy $(\varepsilon,\delta)$-QDP. This claim follows because the estimated lower bound $\mathsf{E}^{\mathrm{LB}}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)>\delta$ implies that $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)>\delta$.
@@
To verify $(\varepsilon,\delta)$-QDP, it is required to compute whether $\widehat{f}(\rho,\sigma,\mathcal{A},\varepsilon)\leq\frac{1}{2}\left(1+\frac{2\delta+e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)$ for all neighboring pairs $(\rho,\sigma)\in\mathcal{Q}$. However, checking this requirement has increasing computational complexity as the cardinality of the set $\mathcal{Q}$ increases. If the privacy requirements can be relaxed so as to encode domain knowledge as in QPP framework, the effective set of $\mathcal{Q}$ may be a small set in some applications of interest.

### v2

@@ VI-B Analysis of Depolarization Mechanism
We now instantiate $\mathcal{A}$ as the depolarizing channel with parameter $p$, denoted as $\mathcal{A}^{p}_{\mathrm{Dep}}$ (as defined in (114)), and proceed to analyze $\mathcal{U}(\mathcal{A}^{p}_{\mathrm{Dep}})$.
@@
$\mathsf{U}(\mathcal{A}_{\mathrm{Dep}}^{p})=1-\frac{p(d^{2}-1)}{d^{2}}\geq\gamma.$ (155)
@@
The proof below relies on observing that the optimization term (in the utility metric) is minimized by setting $\mathcal{B}=\mathcal{I}$, and then evaluating $\left\|\mathcal{I}-\mathcal{A}_{\mathrm{Dep}}^{p}\right\|_{\diamond}$ using the Choi states of the channels $\mathcal{A}_{\mathrm{Dep}}^{p}$ and $\mathcal{I}$, due to the joint covariance of the two channels under unitaries (i.e., $\mathcal{A}_{\mathrm{Dep}}^{p}\circ\mathcal{U}=\mathcal{U}\circ\mathcal{A}_{\mathrm{Dep}}^{p}$ and $\mathcal{I}\circ\mathcal{U}=\mathcal{U}\circ\mathcal{I}$ for every unitary channel $\mathcal{U}$).
@@
where: (a) follows from the unitary invariance of the diamond norm with $\mathcal{U}$ representing a unitary channel; (b) from the commutative property of $\mathcal{A}_{\mathrm{Dep}}^{p}$ with every unitary channel; (c) with $\,\mathsf{d}\mathcal{U}$ denoting the Haar measure over the unitary group and from the left-hand side being independent of $\mathcal{U}$; and (d) from the convexity of the diamond norm.
Next, observe that $\mathcal{B}^{\star}\coloneqq\int\,\mathsf{d}\mathcal{U}\ \mathcal{U}\circ\mathcal{B}\circ\mathcal{U}^{\dagger}$ is a quantum channel, and it is in fact equal to a depolarization channel [65]. Then, $\mathcal{B}^{\star}=\mathcal{A}_{\mathrm{Dep}}^{q}$ for some $q\in[0,1]$.
@@
With the property of joint covariance of $\mathcal{I}$ and $\mathcal{A}_{\mathrm{Dep}}^{p}$ under unitaries [37, Proposition 7.82], we simplify this to
@@
Next, we focus on understanding the privacy-utility tradeoff with respect to the parameter $p$ governing a depolarization mechanism. From Proposition 7, to achieve $\gamma$-utility, we require that
$p\leq\frac{(1-\gamma)d^{2}}{(d^{2}-1)}.$ (168)
Conversely, to achieve $\varepsilon$-QPP in the chosen privacy framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$, from Theorem 3,it suffices for $p$ to satisfy
@@
These two inequalities provide insights into the privacy-utility tradeoff associated with the depolarization mechanism. Consequently, it is essential to carefully tune the parameter $p$ based on the desired utility, characterized by $\gamma$, as well as the privacy parameter $\varepsilon$.
Figure 6: (a) For fixed $d=2$, the figure depicts the optimum utility $\gamma$ for $\varepsilon$ achievable with the depolarization mechanism in Theorem 3. The value of $K$ encodes the domain knowledge available, where $K=1$ corresponds to no such additional information being available. (b) For fixed $K=1$, the figure depicts the optimum utility $\gamma$ for $\varepsilon$ achievable with the depolarization mechanism in Theorem 3 for $d\in\{2,4,8,16\}$.
Effect of domain knowledge: Fig. 6(a) illustrates the optimal utility achievable using the $\varepsilon$-QPP depolarization mechanism presented in Theorem 3. Notably, as the value of $K$ reduces, the attainable utility region expands. The parameter $K$ derived from Theorem 3 represents the domain knowledge accessible and incorporated into the privacy model of the $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ QPP framework. This observation underscores the significance of incorporating such domain knowledge to enhance utility gains while simultaneously ensuring the necessary privacy assurances.
Effect of dimension: In Fig. 6(b), we observe a prominent privacy-utility tradeoff as the dimension increases for the depolarization mechanism presented in Theorem 3. Regarding the utility of the depolarization mechanism (given by $1-\frac{p(d^{2}-1)}{d^{2}}$), we can always establish the following lower bound for every $d$:
$1-\frac{p(d^{2}-1)}{d^{2}}\geq 1-p,$ (170)
where this lower bound is attained as $d\to\infty$. However, the achievable privacy level $\varepsilon$ in (117) degrades at most by an order of $\ln(d)$. Hence, it is crucial to identify the optimal privacy parameters achieved by private mechanisms, particularly in high-dimensional scenarios.
Remark 13 (Application specific privacy-utility tradeoffs).
In the previous analysis concerning the depolarization mechanism in Fig. 4, we chose $\mathcal{E}=\mathcal{I}$, the identity channel. However, it would be an interesting future work to explore the utility for user-specific $\mathcal{E}$ channels. Specifically, we can choose $1-\frac{1}{2}\inf_{\mathcal{B}\in\mathrm{CPTP}}\|\mathcal{E}-\mathcal{B}\circ\mathcal{A}^{p}_{\mathrm{Dep}}\circ\mathcal{E}\|_{\diamond}$ as the utility metric. If $\mathcal{E}$ possesses certain symmetries, one can potentially utilize arguments akin to those presented in the proof of Proposition 7. This investigation could shed light on tailoring privacy mechanisms to specific application needs, leading to more effective privacy-utility tradeoffs.
@@ VII Auditing Privacy Frameworks
Auditing for privacy aims to detect violations in privacy guarantees and reject incorrect algorithms (see [66, 67, 68, 10] for classical approaches). In this section, our focus is on utilizing quantum information theory tools and quantum algorithms to audit the privacy of quantum systems. Specifically, we concentrate on auditing algorithms for QDP guarantees, and it should be noted that these ideas can be extended to audit algorithms for privacy guarantees demanded by QPP (see Remark 16).
The main idea behind auditing classical privacy frameworks (DP and PP) involves translating the privacy requirement into a weaker privacy notion that can be efficiently computed. For example, in [10], sliced mutual information based DP is used to audit for DP. By doing so, algorithms failing to meet the privacy conditions imposed by the relaxed privacy notion are concluded to violate the original privacy requirement. However, a pitfall of this approach is the inability to quantify the gap between the constraints stemming from the original DP or PP notion and the relaxed privacy notions. In other words, even if we verify that the relaxed privacy notion is satisfied, we cannot determine whether the original privacy requirement is also satisfied. In contrast, in this work, we focus on QDP without translating it into a relaxed privacy notion.
@@ VII-A Techniques for Auditing QDP
Using semi-definite programs:
By leveraging the equivalent formulation from Proposition 1 and adopting the specific choices of $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ as provided in Section III-C1, for $(\varepsilon,\delta)$-QDP, we have that
@@
Then, we can compute the left-hand side above by using the SDP formulation of $\overline{\mathsf{D}}^{\delta}\!\left(\cdot\|\cdot\right)$ presented in Lemma 1. This approach is particularly beneficial for low-dimensional setups as the time complexity of SDP computation is polynomial in the dimension of the quantum states. However, it is essential to note that the complexity of this approach grows exponentially with the number of qubits, making it less feasible for higher-dimensional systems.
Additionally, using the equivalent formulation in Remark 6, consider that $(\varepsilon,\delta)$-QDP is equivalent to
@@
with $\gamma\geq 1$ for quantum states $\rho$ and $\sigma$ [15, Eq. (II.2)]. As shown in (45) and (47), the quantity on the right-hand side of (173) can be evaluated by means of an SDP. Then, auditing QDP reduces to computing $\mathsf{E}_{\gamma}\big{(}\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\big{)}$ for $\rho\sim\sigma$. However, similar to the previous approach, the time complexity of this SDP also grows exponentially with the number of qubits. Thus, computing these SDPs remains challenging for higher-dimensional quantum systems.
Using quantum circuits:
Another approach is to borrow the results of [69, 70] and use the connection of $\mathsf{E}_{\gamma}(\rho\|\sigma)$ to the trace distance given in (174). Despite this connection, evaluating ${\mathsf{E}}_{\gamma}(\rho\|\sigma)$ remains computationally challenging, even for quantum computers [69, 70]. Nevertheless, there are proposals for evaluating the trace distance using variational quantum algorithms [71, 72] (which however do not give particular runtimes), and for cases in which the quantum states have low rank [73].
In the subsequent analysis, we explore an approach from [71, 72] (via variational algorithms with parameterized quantum circuits) to estimate the quantity $\left\|\mathcal{A}(\rho)-e^{\varepsilon}\mathcal{A}(\sigma)\right\|_{1}$. Using such an estimate, for a fixed value of the privacy parameter $\varepsilon$, we can validate on which values of $\delta$ the needed guarantees are satisfied.
@@
Firstly, let us focus on how to estimate $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)$ for a fixed $\rho\sim\sigma$ and $\varepsilon$. With the ideas developed in [71, Algorithm 14], we discuss how a process similar to quantum interactive proof can be used for estimating the privacy level. For that, refer to the quantum circuit in Fig. 7: the unitaries $U^{\rho}$ and $U^{\sigma}$ are used to prepare the states $\rho$ and $\sigma$, by tracing out systems $R_{1}$ and $R_{2}$, respectively. Then the algorithm $\mathcal{A}$ is applied on the systems $S_{i}$ for $i\in\{1,2\}$. The unitary $Q$ takes inputs $F_{i},B_{i}$ and outputs $F_{i}^{\prime},T_{i}$, where $F_{i}$ and $T_{i}$ are qubit systems. Finally, both of the $T_{i}$ systems are measured in the standard basis $\{|0\rangle,|1\rangle\}$, and the (classical) output random variable is denoted as $X_{\rho}$ for $i=1$ and $X_{\sigma}$ for $i=2$. Here $X_{\rho}$ and $X_{\sigma}$ take values in $\{0,1\}$. This procedure is repeated a sufficient number of times, and we use the results to estimate $\mathbb{P}(X_{\rho}=0)$ and $\mathbb{P}(X_{\sigma}=1)$.
@@
In quantum complexity terminology, this action could be conducted by a quantum prover who has unbounded computational power (we discuss how to relax this assumption in Remark 14). From [71, Eq. (128)] and the discussion therein, we conclude that
@@
As the number of qubits increases, classical methods (e.g., using SDPs) often become intractable due to the exponential growth in computational complexity. So in light of this, the above approach is desired with the increasing dimension of the quantum system.
However, finding the $Q$ that achieves the optimum utility in (175) is practically infeasible. To relax this assumption, we replace the action of the prover (who finds this $Q$) with a parameterized circuit $Q_{\theta}$. Then we can use (175) as a utility function for training a variational quantum algorithm [74, 75] to estimate a lower bound for (176). With that we obtain a lower bound on $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)$, which we denote as $\mathsf{E}^{\mathrm{LB}}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)$. A lower bound gives a sufficient condition for ruling out algorithms that do not satisfy $(\varepsilon,\delta)$-QDP. This claim follows because the estimated lower bound $\mathsf{E}^{\mathrm{LB}}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)>\delta$ implies that $\mathsf{E}_{e^{\varepsilon}}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)>\delta$.
@@
To verify $(\varepsilon,\delta)$-QDP, it is required to compute whether $\widehat{f}(\rho,\sigma,\mathcal{A},\varepsilon)\leq\frac{1}{2}\left(1+\frac{2\delta+e^{\varepsilon}-1}{e^{\varepsilon}+1}\right)$ for all neighboring pairs $(\rho,\sigma)\in\mathcal{Q}$. However, checking this requirement has increasing computational complexity as the cardinality of the set $\mathcal{Q}$ increases. If the privacy requirements can be relaxed so as to encode domain knowledge as in QPP framework, the effective set $\mathcal{Q}$ may be a small set in some applications of interest. For an example, consider an application where hypothesis testing is carried out between the states $\rho$ and $\sigma$ under privacy constraints where $\mathcal{Q}=\{(\rho,\sigma),(\sigma,\rho)\}$.
Remark 16 (Auditing QPP [^7] Note that the following ideas can also be extended to auditing for variants of QPP in Section IX as well (See also Remark 21).).
To audit for $(\varepsilon,\delta)$-QPP in the framework $(\mathcal{S},\mathcal{Q},\Theta,\bar{\mathcal{M}})$, one can use the ideas described above by choosing $\rho^{\mathcal{R}}$ and $\rho^{\mathcal{T}}$ for all $P_{X}\in\Theta,(\mathcal{R},\mathcal{T})\in\mathcal{Q}$ instead of $\rho$ and $\sigma$. In that case, the complexity of the approach relies on the set of $\mathcal{Q}$ as well as distributions contained in $\Theta$.